A quarterback on a football team has a pass completion rate of 0.62. If, in a given game, he attempts 16 passes, what is the probability that he will complete
(a) 12 passes?
(b) More than half of his passes?
(c) Interpret your result.
(d) Out of the 16 passes, what is the expected number of completions?
Thank you

Answer by ewatrrr(24785)


Hi
Binomial Distribution:  p(pass completion) = .62
Using TI or similarly an inexpensive calculator like a Casio fx-115 ES plus 
n = 16
(a) P(x = 12) = binompdf(16, .62, 12) = .1224
(b) P(x > 6) = 1 - P(x ≤  6)= 1 - binomcdf(16, .62, 6) = 1-.0411 = .9589
(c) Expected = .62 * 16 = 9.92  or 9 (rounded down to whole number)

Important You are  comfortable using Your Calculator.
Wish You the Best in your Studies.